Hi
Can someone please help with question
If a room has a width of 5 m, a length of 6 m,and a height of 5 m. How many tiles do you need to decorate your room if each tile is 6 cm in length and width? Note that the ceiling and the floor do not need the tiles. If each tile cost $0.5, how much money will you need to spend on decorating your room?
Thanks in advance

Found 2 solutions by addingup, jim_thompson5910:
Answer by addingup(3677)

1 meter = 100 cm
5 x 100 = 500
6 x 100 = 600
- - - - - - - - - - - - -
500 x 500 = 25000; 25000 x 2 = 50000 2 walls measure 50000 square cm.
600 x 500 = 30000; 30000 x 2 = 60000 2 walls measure 60000 square cm.
- - - - - - - - - - - - -
50000 + 60000 = 110000 cm^2 total wall surface to be tiled
6^2 = 36 cm^2 each tile
110000/36 = 3055.556 tiles, which means you need 3056 tiles and you'll have a small piece left over (0.444 of a tile exactly)

Answer by jim_thompson5910(35256)


The tutor @addingup has the right idea, but made a few typos

They wrote 500 x 500 = 25000
when it should be 500 x 500 = 250000 (note the extra zero)

Similarly,
600 x 500 = 30000
should be
600 x 500 = 300000

The total area of all four walls is
300000+300000+250000+250000 = 1,100,000 square cm

Divide this over the area of 1 tile (36 cm^2) and we get
1,100,000/36 = 30,555.5555555556
which rounds up to 30,556

You need 30,556 tiles
At $0.50 per tile, the total cost is
30,556*0.5 = 15,278 dollars

This is one expensive project. The reason why is because of the size of each wall being fairly massive, compared to the relatively very small tile size. That leads to a lot of tiles needed.